Do I have this rate of change problem correct?
f(x)=4-x^2; x=9, x=9+h
f(9+h)-f(9)/(9+h-9) = 4-(9+h)-4-(9)^2/h
=-162+h/h = -162 because the h's cancel out? Not sure if I did this correctly.
Please explain, thank you!

Answer by stanbon(75887)

f(x)=4-x^2; x=9, x=9+h
------------------------------
f(9+h) = 4 - (9+h)^2 = 4 -81 -18h -h^2
f(9) = 4- (81) = -77
----
Rate of change at x = 9 is [f(9+h)-f(9)]/h
rate of change at x = 9 is [-18h-h^2]/h = -18 -h
Note:: As h goes to zero, rate goes to -18
----------
Cheers,
Stan H.